# Editable Concept Bottleneck Models

### v1

@@ 4.3 Data-level Editable CBM
For dataset $\mathcal{D}=\{(x_{i},y_{i},c_{i})\}_{i=1}^{n}$, given a set of data $z_{r}=(x_{r},y_{r},c_{r})$, $r\in G$ to be removed. Suppose the updated concept predictor $\hat{g}_{-z_{G}}$ is defined by (9), then we have the following approximation for $\hat{g}_{-z_{G}}$
@@
Compared with the original loss before unlearning in (2), we can observe two changes in (11). First, we remove $|G|$ data points in the loss function $L_{Y}$. Secondly, the input for the loss is also changed from $\hat{g}(x_{i})$ to $\hat{g}_{-z_{G}}$. Therefore, it is difficult to estimate directly with an influence function. Here we introduce an intermediate label predictor as
@@
For dataset $\mathcal{D}=\{(x_{i},y_{i},c_{i})\}_{i=1}^{n}$, given a set of data $z_{r}=(x_{r},y_{r},c_{r})$, $r\in G$ to be removed. The intermediate label predictor $\tilde{f}_{-z_{G}}$ is defined in (12). Then we have
@@
Acceleration via EK-FAC. As we mentioned in Section 3, as the loss function in CBMs is non-convex, the Hessian matrices in all our theorems may not be well-defined. We can use the EK-FAC approach, i.e., using $\hat{H}=G+\lambda I$ to approximate the Hessian, where $G$ is the Fisher information matrix, and $\lambda$ is a small damping term used to ensure the positive definiteness. See Appendix A.1 for using EK-FAC to CBMs. Also, see Algorithm E-E in the Appendix for the detailed EK-FAC-based algorithms for our three levels, whose original (Hessian) versions are in Algorithm E-E, respectively.
@@ 5 Experiments
In this section, we give our main experimental results on utility evaluation, edition efficiency, and interpretability evaluation. Details and additional results are in Appendix F due to space limit.
@@ 5.1 Experimental Settings
Dataset. We utilize two datasets: X-ray grading (OAI) nevitt2006osteoarthritis and Bird identification (CUB) wah2011caltech . X-ray grading (OAI) is a multi-center observational study of knee osteoarthritis, which comprises 36,369 data points. Specifically, we configure n=10 concepts that characterize crucial osteoarthritis indicators such as joint space narrowing, osteophytes, and calcification. Bird identification (CUB) consists of 11,788 data points, which belong to 200 classes and include 112 binary attributes to describe detailed visual features of birds. We follow the same network architecture and settings outlined in koh2020concept .
Baselines. We employ three baseline methods: Retrain, CBM-IF, and ECBM. Retraining: We will retrain the CBM from scratch by removing the samples, concept labels, or concepts. CBM-IF: This method is a direct implementation of our previous Theorems of model updates for the three settings, see Algorithms E-E for details. ECBM: As we discussed above, all of our model updates can be further accelerated via EK-FAC, ECBM corresponds to the EK-FAC accelerated version of Algorithms E-E (refer to Algorithms E-E).
@@ 5.2 Evaluation of Utility and Editing Efficiency
Our experimental results, as illustrated in Table 1, demonstrate the effectiveness of ECBMs compared to traditional retraining and CBM-IF, particularly emphasizing computational efficiency without compromising accuracy. Specifically, ECBMs achieved an F1 score close to that of retraining (0.8808 vs. 0.8825) while significantly reducing the runtime from 31.44 seconds to 8.29 seconds. This pattern was consistent in the CUB dataset, where the runtime was decreased from 27.88 seconds for retraining to 7.03 seconds for ECBMs, with a negligible difference in the F1 score (0.7971 to 0.7963). These results highlight the potential of ECBMs to provide substantial time savings—approximately 22-30% of the computational time required for retraining—while maintaining comparable accuracy. Compared to CBM-IF, ECBM also showed a slight reduction in runtime and a significant improvement in F1 score. The former verifies the effective acceleration of our algorithm by EK-FAC. This efficiency is particularly crucial in scenarios where frequent updates to model annotations are needed, confirming the utility of ECBMs in dynamic environments where speed and accuracy are critical.
@@
Table 1: Performance comparison of different methods on the OAI and CUB datasets.
Edit Level Method OAI CUB
F1 score RT (second) F1 score RT (second)
Concept Label Retrain 0.8825$\pm$0.0054 31.44 0.7971$\pm$0.0066 27.88
CBM-IF(Ours) 0.8639$\pm$0.0033 16.31 0.7699$\pm$0.0035 14.39
ECBM(Ours) 0.8808$\pm$0.0039 8.29 0.7963$\pm$0.0050 7.03
Concept Retrain 0.8448$\pm$0.0191 27.33 0.7811$\pm$0.0047 28.41
CBM-IF(Ours) 0.8214$\pm$0.0071 17.38 0.7579$\pm$0.0065 15.70
ECBM(Ours) 0.8403$\pm$0.0090 8.30 0.7787$\pm$0.0058 6.43
Data Retrain 0.8811$\pm$0.0065 33.72 0.7838$\pm$0.0051 28.08
CBM-IF(Ours) 0.8472$\pm$0.0046 17.84 0.7623$\pm$0.0031 15.86
ECBM(Ours) 0.8797$\pm$0.0038 8.81 0.7827$\pm$0.0088 7.11
@@ 5.3 Results on Interpretability
Influence function in ECBM can measure the importance of concepts. As the original motivation of influence function is to calculate the importance score of each sample. Here we will show that the influence function for the concept level in Theorem 4.3 can be used to calculate the importance of each concept in CBMs, which provides an explainable tool for CBMs. In detail, we first select 1-10 most influential and 1-10 least influential concepts by our influence function. Then we will remove these concepts and update the model via retraining or our ECBM, and analyze the change (F1 Score Difference) w.r.t. the original CBM before removal.
The results in Figure 3(a) demonstrate that when we remove the $1$-$10$ most influential concepts identified by the ECBM method, the F1 score decreases by more than $0.025$ compared to the CBM before removal. In contrast, Figure 3(b) shows that the change in the F1 score remains consistently below $0.005$ when removing the least influential concepts. These findings strongly indicate that the influence function in ECBM can successfully determine the importance of concepts. Furthermore, we observe that the gap between the F1 score of retraining and ECBM is consistently smaller than $0.005$, and even smaller in the case of least important concepts. This further suggests that when ECBM edits various concepts, its performance is very close to the ground truth.
@@
Visualization. Since CBM is an explainable model, we aim to evaluate the interpretability of our ECBM (compared to the retraining). We will present some visualization results for the concept-level edit. Figure 4 presents the top $10$ most influential concepts and their corresponding predicted concept labels obtained by our ECBM and the retrain method after randomly deleting concepts. Our ECBM can provide explanations for which concepts are crucial and how they determine the predicted label. Specifically, among the top $10$ most important concepts in the ground truth (retraining), ECBM can accurately recognize $9$ with their correct concept labels. For instance, we correctly identify "has_upperparts_color::orange", "has_upper_tail_color::red", and "has_breast_color::black" as some of the most important concepts when predicting categories. Additional visualization results Figure 5 and 6 are in the appendix.
@@ 6 Conclusion
Appendix A Acceleration for Influence Function
@@
Firstly, assume that the derivatives of the weights in different layers are uncorrelated, which implies that $G_{\hat{\theta}}$ has a block-diagonal structure. Suupose $\hat{g}_{\theta}$ can be denoted by $\hat{g}_{\theta}(x)=g_{\theta_{L}}\circ\cdots\circ g_{\theta_{l}}\circ\cdots\circ g_{\theta_{1}}(x)$ where $l\in[L]$. We fold the bias into the weights and vectorize the parameters in the $l$-th layer into a vector $\theta_{l}\in\mathbb{R}^{d_{l}}$, $d_{l}\in\mathbb{N}$ is the number of $l$-th layer parameters. Then $G_{\hat{\theta}}$ can be reaplcaed by $\left(G_{1}(\hat{\theta}),\cdots,G_{L}(\hat{\theta})\right)$, where $G_{l}(\hat{\theta})\triangleq n^{-1}\sum_{i=1}^{n}\nabla_{\hat{\theta}_{l}}\ell_{i}\nabla_{\theta_{l}}\ell_{i}^{\mathrm{T}}$. Denote $h_{l}$, $o_{l}$ as the output and pre-activated output of $l$-th layer. Then $G_{l}(\theta)$ can be approximated by
@@
Furthermore, in order to accelerate transpose operation and introduce the damping term, perform eigenvalue decomposition of matrix ${\Omega}_{l-1}$ and ${\Gamma_{l}}$ and obtain $Q_{\Omega}{\Lambda}_{{\Omega}}{Q}_{{\Omega}}^{\top}$ and ${Q}_{\Gamma}{\Lambda}_{\Gamma}{Q}_{\Gamma}^{\top}$. Then the inverse of $\hat{H}_{l}(\theta)$ can be obtained by
@@
Besides, [14] proposed a new method that corrects the error in equation 14 which sets the $i$-th diagonal element of $\Lambda_{\Omega_{l-1}}\otimes\Lambda_{\Gamma_{l}}$ as $\Lambda_{ii}^{*}=n^{-1}\sum_{j=1}^{n}\left(\left(Q_{\Omega_{l-1}}\otimes Q_{\Gamma_{l}}\right)\nabla_{\theta_{l}}\ell_{j}\right)^{2}_{i}.$
A.1 EK-FAC for CBMs

### v2

@@ 4.3 Data-level Editable CBM
For dataset $\mathcal{D}=\{(x_{i},y_{i},c_{i})\}_{i=1}^{n}$, given a set of data $z_{r}=(x_{r},y_{r},c_{r})$, $r\in G$ to be removed. Suppose the updated concept predictor $\hat{g}_{-z_{G}}$ is defined by equation 9, then we have the following approximation for $\hat{g}_{-z_{G}}$
@@
Compared with the original loss before unlearning in equation 2, we can observe two changes in equation 11. First, we remove $|G|$ data points in the loss function $L_{Y}$. Secondly, the input for the loss is also changed from $\hat{g}(x_{i})$ to $\hat{g}_{-z_{G}}$. Therefore, it is difficult to estimate directly with an influence function. Here we introduce an intermediate label predictor as
@@
For dataset $\mathcal{D}=\{(x_{i},y_{i},c_{i})\}_{i=1}^{n}$, given a set of data $z_{r}=(x_{r},y_{r},c_{r})$, $r\in G$ to be removed. The intermediate label predictor $\tilde{f}_{-z_{G}}$ is defined in equation 12. Then we have
@@
Acceleration via EK-FAC. As we mentioned in Section 3, as the loss function in CBMs is non-convex, the Hessian matrices in all our theorems may not be well-defined. We can use the EK-FAC approach, i.e., using $\hat{H}=G+\lambda I$ to approximate the Hessian, where $G$ is the Fisher information matrix, and $\lambda$ is a small damping term used to ensure the positive definiteness. See Appendix B.1 for using EK-FAC to CBMs. Also, see Algorithm F-F in the Appendix for the detailed EK-FAC-based algorithms for our three levels, whose original (Hessian) versions are in Algorithm F-F, respectively.
@@ 5 Experiments
In this section, we demonstrate our main experimental results on utility evaluation, edition efficiency, and interpretability evaluation. Details and additional results are in Appendix G due to space limit.
@@ 5.1 Experimental Settings
Dataset. We utilize three datasets: X-ray grading (OAI) (Nevitt et al., 2006), Bird identification (CUB) (Wah et al., 2011) and Large-scale CelebFaces Attributes dataset (CelebA) (Liu et al., 2015). OAI is a multi-center observational study of knee osteoarthritis, which comprises 36,369 data points. Specifically, we configure $n=10$ concepts that characterize crucial osteoarthritis indicators such as joint space narrowing, osteophytes, and calcification. Bird identification (CUB) consists of 11,788 data points, which belong to 200 classes and include 112 binary attributes to describe detailed visual features of birds. CelebA comprises over 200,000 celebrity images, each annotated with 40 binary attributes that detail facial features, such as hair color, eyeglasses, and smiling. As the dataset lacks predefined classification tasks, following Espinosa Zarlenga et al. (2022), we designate 8 attributes as labels and the remaining 32 attributes as concepts. For all the above datasets, we follow the same network architecture and settings outlined in Koh et al. (2020).
Baselines. We employ three baseline methods: Retrain, CBM-IF, and ECBM. Retrain: We retrain the CBM from scratch by removing the samples, concept labels, or concepts from the training set. CBM-IF: This method is a direct implementation of our previous theorems of model updates in the three settings. See Algorithms F-F in Appendix for details. ECBM: As we discussed above, all of our model updates can be further accelerated via EK-FAC, ECBM corresponds to the EK-FAC accelerated version of Algorithms F-F (refer to Algorithms F-F in Appendix).
@@ 5.2 Evaluation of Utility and Editing Efficiency
Our experimental results, as illustrated in Table 1, demonstrate the effectiveness of ECBMs compared to traditional retraining and CBM-IF, particularly emphasizing computational efficiency without compromising accuracy. Specifically, ECBMs achieved F1 scores close to those of retraining (0.8808 vs. 0.8825) while significantly reducing the runtime from 31.44 seconds to 8.29 seconds. This pattern is consistent in the CUB dataset, where the runtime was decreased from 27.88 seconds for retraining to 7.03 seconds for ECBMs, with a negligible difference in the F1 score (0.7971 to 0.7963). These results highlight the potential of ECBMs to provide substantial time savings—approximately 22-30% of the computational time required for retraining—while maintaining comparable accuracy. Compared to CBM-IF, ECBM also showed a slight reduction in runtime and a significant improvement in F1 score. The former verifies the effective acceleration of our algorithm by EK-FAC. This efficiency is particularly crucial in scenarios where frequent updates to model annotations are needed, confirming the utility of ECBMs in dynamic environments where running time and accuracy are critical.
@@
Table 1: Performance comparison of different methods on the three datasets.
Edit Level Method OAI CUB CelebA
F1 score RT (second) F1 score RT (second) F1 score RT (second)
Concept Label Retrain 0.8825$\pm$0.0054 31.44 0.7971$\pm$0.0066 27.88 0.3827$\pm$0.0272 57.60
CBM-IF(Ours) 0.8639$\pm$0.0033 16.31 0.7699$\pm$0.0035 14.39 0.3561$\pm$0.0134 34.93
ECBM(Ours) 0.8808$\pm$0.0039 8.29 0.7963$\pm$0.0050 7.03 0.3845$\pm$0.0327 15.67
Concept Retrain 0.8448$\pm$0.0191 27.33 0.7811$\pm$0.0047 28.41 0.3776$\pm$0.0350 68.94
CBM-IF(Ours) 0.8214$\pm$0.0071 17.38 0.7579$\pm$0.0065 15.70 0.3609$\pm$0.0202 35.56
ECBM(Ours) 0.8403$\pm$0.0090 8.30 0.7787$\pm$0.0058 6.43 0.3761$\pm$0.0280 15.99
Data Retrain 0.8811$\pm$0.0065 33.72 0.7838$\pm$0.0051 28.08 0.3797$\pm$0.0375 65.60
CBM-IF(Ours) 0.8472$\pm$0.0046 17.84 0.7623$\pm$0.0031 15.86 0.3536$\pm$0.0166 40.08
ECBM(Ours) 0.8797$\pm$0.0038 8.81 0.7827$\pm$0.0088 7.11 0.3748$\pm$0.0347 16.75
@@ 5.3 Results on Interpretability
Influence function in ECBM can measure the importance of concepts. The original motivation of the influence function is to calculate the importance score of each sample. Here, we will show that the influence function for the concept level in Theorem 4.3 can be used to calculate the importance of each concept in CBMs, which provides an explainable tool for CBMs. In detail, we conduct our experiments on the CUB dataset. We first select 1-10 most influential and 1-10 least influential concepts by our influence function. Then, we will remove these concepts and update the model via retraining or our ECBM and analyze the change (F1 Score Difference) w.r.t. the original CBM before removal.
The results in Figure 3(a) demonstrate that when we remove the 11-1010 most influential concepts identified by the ECBM method, the F1 score decreases by more than 0.0250.025 compared to the CBM before removal. In contrast, Figure 3(b) shows that the change in the F1 score remains consistently below $0.005$ when removing the least influential concepts. These findings strongly indicate that the influence function in ECBM can successfully determine the importance of concepts. Furthermore, we observe that the gap between the F1 score of retraining and ECBM is consistently smaller than $0.005$, and even smaller in the case of least important concepts. This further suggests that when ECBM edits various concepts, its performance is very close to the ground truth.
@@
ECBMs can erase data influence. For the data level, ECBMs aim to facilitate an efficient removal of samples. We perform membership inference attacks (MIAs) to provide direct evidence that ECBMs can indeed erase data influence. MIA is a privacy attack that aims to infer whether a specific data sample was part of the training dataset used to train a model. The attacker exploits the model’s behavior, such as overconfidence or overfitting, to distinguish between training (member) and non-training (non-member) data points. In MIAs, the attacker typically queries the model with a data sample and observes its prediction confidence or loss values, which tend to be higher for members of the training set than non-members (Shokri et al., 2017).
To quantify the success of these edits, we calculate the RMIA (Removed Membership Inference Attack) score for each category. The RMIA score is defined as the model’s confidence in classifying whether a given sample belongs to the training set. Lower RMIA values indicate that the sample behaves more like a test set (non-member) sample Zarifzadeh et al. (2024). This metric is especially crucial for edited samples, as a successful ECBM should make the removed members behave similarly to non-members, reducing their membership vulnerability. See Appendix G for its definition.
We conducted experiments by randomly selecting 200 samples from the training set (members) and 200 samples from the test set (non-members) of the CUB dataset. We calculated the RMIA scores for these samples and plotted their frequency distributions, as shown in Figure 4(a). The mean RMIA score for non-members was 0.049465, while members had a mean score of 0.063505. Subsequently, we applied ECBMs to remove the 200 training samples from the model, updated the model parameters, and then recalculated the RMIA scores. After editing, the mean RMIA score for the removed-members decreased to 0.052105, significantly closer to the non-members’ mean score. This shift in RMIA values demonstrates the effectiveness of ECBMs in editing the model, as the removed members now exhibit behavior closer to that of non-members. The post-editing RMIA score distributions are shown in Figure 4(b). These results provide evidence of the effectiveness of ECBMs in editing the model’s knowledge about specific samples.
(a) RMIA Score Before Editing
(b) RMIA Score After Editing
Figure 4: RMIA scores of data before and after removal.
Visualization. Since CBM is an explainable model, we aim to evaluate the interpretability of our ECBM (compared to the retraining). We will present some visualization results for the concept-level edit. Figure 5 presents the top 1010 most influential concepts and their corresponding predicted concept labels obtained by our ECBM and the retrain method after randomly deleting concepts for the CUB dataset. Our ECBM can provide explanations for which concepts are crucial and how they determine the predicted label. Specifically, among the top $10$ most important concepts in the ground truth (retraining), ECBM can accurately recognize $9$ with their correct concept labels. For instance, we correctly identify "has_upperparts_color::orange", "has_upper_tail_color::red", and "has_breast_color::black" as some of the most important concepts when predicting categories. Additional visualization results under data level and concept-label level on OAI and CUB datasets are included in Appendix G.2.
@@ 6 Conclusion
Appendix A Influence Function
Consider a neural network $\hat{\theta}=\arg\min_{\theta}\sum_{i=1}^{n}\ell(z_{i},\theta)$ with loss function $L$ and dataset $D=\{z_{i}\}_{i=1}^{n}$. That is $\hat{\theta}$ minimize the empirical risk
$R(\theta)=\sum_{i=1}^{n}L(z_{i},\theta)$
Assume $R$ is strongly convex in $\theta$. Then $\theta$ is uniquely defined. If we remove a point $z_{m}$ from the training dataset, the parameters become $\hat{\theta}_{-z_{m}}=\arg\min_{\theta}\sum_{i\neq m}L(z_{i},\theta)$. Up-weighting $z_{m}$ by $\epsilon$ small enough, then the revised risk ${R(\theta)}^{{}^{\prime}}=\frac{1}{n}\sum_{i=1}^{n}L(z_{i};\theta)+\epsilon L(z_{m};\theta)$ is still strongly convex. Then the response function $\hat{\theta}_{\epsilon,-z_{m}}={R(\theta)}^{{}^{\prime}}$ is also uniquely defined. The parameter change is denoted as $\Delta_{\epsilon}=\hat{\theta}_{\epsilon,-z_{m}}-\hat{\theta}$. Since $\hat{\theta}_{\epsilon,-z_{m}}$ is the minimizer of ${R(\theta)}^{{}^{\prime}}$, we have the first-order optimization condition as
$\nabla_{\hat{\theta}_{\epsilon,-z_{m}}}R(\theta)+\epsilon\cdot\nabla_{\hat{\theta}_{\epsilon,-z_{m}}}L(z_{m},\hat{\theta}_{\epsilon,-z_{m}})=0$
Since $\hat{\theta}_{\epsilon,-z_{m}}\rightarrow\hat{\theta}as\epsilon\rightarrow 0$, we perform a Taylor expansion of the right-hand side:
$\left[\nabla R(\hat{\theta})+\epsilon\nabla L(z_{m},\hat{\theta})\right]+\left[\nabla^{2}R(\hat{\theta})+\epsilon\nabla^{2}L(z_{m},\hat{\theta})\right]\Delta_{\epsilon}\approx 0$
Noting $\epsilon\nabla^{2}L(z_{m},\hat{\theta})\Delta_{\epsilon}$ is $o(\|\Delta_{\epsilon}\|)$ term, which is smaller than other parts, we drop it in the following analysis. Then the Taylor expansion equation becomes
$\left[\nabla R(\hat{\theta})+\epsilon\nabla L(z_{m},\hat{\theta})\right]+\nabla^{2}R(\hat{\theta})\cdot\Delta_{\epsilon}\approx 0$
Solving for $\Delta_{\epsilon}$, we obtain:
$\Delta_{\epsilon}=-\left[\nabla^{2}R(\hat{\theta})+\epsilon\nabla^{2}L(z,\hat{\theta})\right]^{-1}\left[\nabla R(\hat{\theta})+\epsilon\nabla L(z,\hat{\theta})\right].$
Remember $\theta$ minimizes $R$, then $\nabla R(\hat{\theta})=0$. Dropping $o(\epsilon)$ term, we have
$\Delta_{\epsilon}=-\epsilon\nabla^{2}R(\hat{\theta})^{-1}\nabla L(z,\hat{\theta}).$
$\left.\frac{d\hat{\theta}_{\epsilon,-z_{m}}}{d\epsilon}\right|_{\epsilon=0}=\left.\frac{d\Delta_{\epsilon}}{d\epsilon}\right|_{\epsilon=0}=-H_{\hat{\theta}}^{-1}\nabla L(z,\hat{\theta})\equiv\mathcal{I}_{up,params}(z).$
Besides, we can obtain the approximation of $\hat{\theta}_{-z_{m}}$ directly by $\hat{\theta}_{-z_{m}}\approx\hat{\theta}+\mathcal{I}_{up,params}(z)$.
Appendix B Acceleration for Influence Function
@@
Firstly, assume that the derivatives of the weights in different layers are uncorrelated, which implies that $G_{\hat{\theta}}$ has a block-diagonal structure. Suppose $\hat{g}_{\theta}$ can be denoted by $\hat{g}_{\theta}(x)=g_{\theta_{L}}\circ\cdots\circ g_{\theta_{l}}\circ\cdots\circ g_{\theta_{1}}(x)$ where $l\in[L]$. We fold the bias into the weights and vectorize the parameters in the $l$-th layer into a vector $\theta_{l}\in\mathbb{R}^{d_{l}}$, $d_{l}\in\mathbb{N}$ is the number of $l$-th layer parameters. Then $G_{\hat{\theta}}$ can be reaplcaed by $\left(G_{1}(\hat{\theta}),\cdots,G_{L}(\hat{\theta})\right)$, where $G_{l}(\hat{\theta})\triangleq n^{-1}\sum_{i=1}^{n}\nabla_{\hat{\theta}_{l}}\ell_{i}\nabla_{\theta_{l}}\ell_{i}^{\mathrm{T}}$. Denote $h_{l}$, $o_{l}$ as the output and pre-activated output of $l$-th layer. Then $G_{l}(\theta)$ can be approximated by
@@
Furthermore, in order to accelerate transpose operation and introduce the damping term, perform eigenvalue decomposition of matrix ${\Omega}_{l-1}$ and ${\Gamma_{l}}$ and obtain the corresponding decomposition results as $Q_{\Omega}{\Lambda}_{{\Omega}}{Q}_{{\Omega}}^{\top}$ and ${Q}_{\Gamma}{\Lambda}_{\Gamma}{Q}_{\Gamma}^{\top}$. Then the inverse of $\hat{H}_{l}(\theta)$ can be obtained by
@@
Besides, George et al. (2018) proposed a new method that corrects the error in equation 14 which sets the $i$-th diagonal element of $\Lambda_{\Omega_{l-1}}\otimes\Lambda_{\Gamma_{l}}$ as $\Lambda_{ii}^{*}=n^{-1}\sum_{j=1}^{n}\left(\left(Q_{\Omega_{l-1}}\otimes Q_{\Gamma_{l}}\right)\nabla_{\theta_{l}}\ell_{j}\right)^{2}_{i}.$
B.1 EK-FAC for CBMs